OpenStudy (anonymous):

Write the following equation in the general form Ax + By + C = 0. 2x + y = 6 2x + y - 6 = 0 2x + y + 6 = 0 2x - y - 6 = 0

OpenStudy (anonymous):

(2)x+(1)y+(-6)=0 (2)x+(1)y+(-6)=0 (2)x+(1)y+6=0 (2)x+(-1)y+(-6)=0